OpenStudy (anonymous):

solve using substitution: 7x-y=83y-5x=8

OpenStudy (anonymous):

7x - y = 83 y - 5x = 8 just checking...are these your two equations ?

OpenStudy (anonymous):

or is it 7x - y = 8 3y - 5x = 8

OpenStudy (anonymous):

its 7x-y+83 and y-5x=8 two equations